determine whether the graphs of the pair of lines are parallel.
x + 6 = y
y - x = -2
What is the slope of the line y + 6 =Y?

Answer by stanbon(75887)

determine whether the graphs of the pair of lines are parallel.
x + 6 = y
slope = 1; y-intercept = 6
------------------------------
y - x = -2
y = x - 2
slope = 1 ; y-intercept = -2
-------------------------------------
The lines are parallel because the slopes
are the same AND the y-intercepts are different.
----------------------------------------------
What is the slope of the line y + 6 =Y?
The equation is a contradiction.
Subtracting y from both sides you get:
6 = 0 which is meaningless.
================================
Cheers,
Stan H.